OpenStudy (anonymous):

Determine if the sequence 5, 30, 180, 1080... is arithmetic, geometric, or niether. A. Arithmetic B. Geometric C. Niether What is the common ratio of this geometric sequence? 1 50 2 30 3 18 4 10.8 5 6.48 A. -20 B. 0.6 C. 1.67 D. 20

OpenStudy (hunus):

Do you know if it's geometric or arithmetic or neither?

OpenStudy (hunus):

If a sequence is geometric, the common ratio will be the number you get when you divide the terms by the ones before it. If these numbers aren't the same for every term, the sequence isn't geometric. For example 1, 2, 4, 8, 16 16/8 = 2 8/4 = 2 4/2 = 2 2/1 = 2 the common ratio is two If it is arithmetic, when you subtract a number from the one after it you will get the common difference For example 1, 3, 5, 7, 9... 9-7=2 7-5=2 5-3=2 3-1=2 The common difference is two If neither of these work it is neither geometric nor arithmetic